OpenStudy (bossimbacon):

What is the missing exponent? (10^-6) ___ = 10 ^-18

OpenStudy (bossimbacon):

@Destinymasha

OpenStudy (bossimbacon):

@Destinymasha plz help

OpenStudy (destinymasha):

Okay, do you know what exponents are , and what they mean?

OpenStudy (bossimbacon):

yeah i do but can you tell me the answer:/ my grade depends on it

OpenStudy (destinymasha):

I can't just tell you the answer >.< You have to work it out :3 10^ -6 = 10 x 10 x 10 x 10 x 10 x 10 (the answer will be flipped around since it's a negative power)

OpenStudy (destinymasha):

You can plug that into a calculator ^.^

OpenStudy (bossimbacon):

1000000

OpenStudy (destinymasha):

Yes, but you have to flip it around since it's a negative. So it would be 0.000001

OpenStudy (bossimbacon):

oh ;o

OpenStudy (bossimbacon):

if its 0.000001 then is that the answer?

OpenStudy (destinymasha):

Err..does this question have any options? I'm not sure what they mean by 'missing exponent' :o

OpenStudy (bossimbacon):

so you don't know the answer :O

OpenStudy (destinymasha):

No, actually, I'm kinda of confused >.< Math isn't my best subject, sorry~ But your problem seems incomplete

OpenStudy (bossimbacon):

What is the value of expression? 6^−7 • 6^4 A. 6−3 B. 6−11 C. 6−28 D. 611

OpenStudy (bossimbacon):

can you tell me this at least?

OpenStudy (destinymasha):

Symbols aren't working >.<

OpenStudy (bossimbacon):

What is the value of expression? 6^-7 • 6^4 A. 6^-3 B. 6^-11 C. 6^-28 D. 6^11

OpenStudy (bossimbacon):

the symbols that are there is a multiplication symbol

OpenStudy (destinymasha):

Okie Gotcha - So..6^-7 x 6^4 = this super long number (0.00462962962) And the only one that matches is...6^-3 , A.

OpenStudy (bossimbacon):

thanks!

OpenStudy (bossimbacon):

can you help me with another one?

OpenStudy (destinymasha):

You're welcome (: And I can try c:

OpenStudy (bossimbacon):

@Destinymasha

OpenStudy (destinymasha):

I'm here, give me a second :P

OpenStudy (destinymasha):

Okay, do you understand the concept at least?

OpenStudy (bossimbacon):

yeah, i do:D i just need answers I'm lazy. ill medal

OpenStudy (bossimbacon):

i would fan but I'm already your fan

OpenStudy (destinymasha):

When it comes to powers, I use calculators :P What they are asking is 3^-6 divided by 3^? = 81 ( 3^4) 3^-6 divided by 3^-10 = 81 (: So your answer would be A

OpenStudy (destinymasha):

And thank you (:

OpenStudy (bossimbacon):

@Destinymasha also can you answer this:D

OpenStudy (destinymasha):

So..you said you knew how to do this stuff, right? <.<

OpenStudy (bossimbacon):

yeah i do haha I'm just in a hrry:p

OpenStudy (destinymasha):

Umm....okie *~* 7^-3 x 7^9 divided by 7^-8 = 678223072849, which is the same as 7^14 :P

OpenStudy (bossimbacon):

thanks! and one more? :P

OpenStudy (bossimbacon):

IM GONNA GIVE YOU ONE BIG FAT MEDEAL D/W

OpenStudy (destinymasha):

Haha

OpenStudy (destinymasha):

Umm...8^-5 divided by (8^-7)^2 = 134217728. Which is equivalent to 8^9

OpenStudy (bossimbacon):

SO THE ANSWER IS ^9 RIGHT

OpenStudy (destinymasha):

Yes

OpenStudy (bossimbacon):

OMG TYSM ILYSM<3 HERES YOUR MEDAL TEACHER :D

OpenStudy (destinymasha):

Haha, thank you :P And you're very welcome (: